Tina and David can produce any three combinations of goods X and Y as shown below:

TINA

DAVID

l60X and 0Y

80X and 0Y

80X and 60Y

40X and 50Y

0X and 120Y

0X and 100Y

  1. Draw a production possibility frontier (curve) for Tina, and one for David.
  2. If Tina is producing 80X and 60Y, what is her opportunity cost for producing one more of X? If David is producing 40X and 50Y, what is his opportunity cost for producing one more unit of X?
  3. Who has a comparative advantage in producing good X? Of good Y? Explain your answer.

d.   BONUS. What is a possible trade agreement that will make both Tina and David better off through specialization and trade (there are several possibilities). Assume that they originally are producing both goods separately. That is, start with Tina producing 80X and 60Y and David producing 40X and 50Y. Then show how much more each can gain by producing only one product and trading for the other, so that both have more of each product.
Answer #1

a.

b.

TINA

DAVID

160X and 0Y

80X and 0Y

80X and 60Y

40X and 50Y

0X and 120Y

0X and 100Y

For Tina:

Tina can produce 80X and 60Y or 160X and 0Y

i.e., In order to increase the production by 160X-80X = 80X, Tina must forego 60Y

Therefore, the opportunity cost of producing 1 more unit of X = (60/80)Y = 0.75Y

For David:

David can produce 40X and 50Y or 80X and 0Y

i.e., In order to increase the production by 80X-40X = 40X, David must forego 50Y

Therefore, the opportunity cost of producing 1 more unit of X = (50/40)Y = 1.20Y

c. Comparative advantage occurs when the opportunity cost of producing a good for one entity is less than the opportunity cost for other entity.

As observed in part (b), the opportunity cost of producing 1 additional unit of X is more for David compared to Tina. Therefore, Tina has a comparative advantage in the production of X

The opportunity cost of producing 1 additional unit of Y for Tina = (80/60)X = 1.33X

The opportunity cost of producing 1 additional unit of Y for David = (40/50)X = 0.80X

The opportunity cost of producing 1 additional unit of Y is more for Tina compared to David. Therefore, David has a comparative advantage in the production of good Y.

d. Trade would be favourable for both Tina and David when the terms of trade are exactly in between their respective opportunity costs

Let us assume that the terms of trade are 1 unit of good X is exchanged for 1 unit of good Y

According to their respective comparative advantages, Tina specializes in the production of X and David specializes in the production of Y

Tina produces 160X and David produces 120Y

Let the exchange be 65 units of X for 65 units of Y

Then, TIna's final consumption = 95X and 65Y (against the initial consumption of 80X and 60Y)

David's final consumption = 65X and 55Y (against the initial consumption of 40X and 50Y)

Therefore, Tina can gain 15X and 5Y while David can gain 25X and 5Y